OpenStudy (anonymous):

For the function y= -2+5sin(pi/12(x-2)), what is the minimum value?

OpenStudy (tkhunny):

Where is \(\sin\left(\dfrac{\pi}{12}(x-2)\right) = -1\) ? Wait, that's WHERE is the minimum value. WHAT is a much easier question. Just look at it and say -2-5 = -7. Where were you stuck?